OpenStudy (anonymous):

If g(x) = 3x - 5 find g-1(x)

OpenStudy (anonymous):

Help Please

OpenStudy (anonymous):

To find the inverse, plug in g-1(x) into g(x) and solve for the x.

OpenStudy (anonymous):

Err solve for the g^-1(x) rather

OpenStudy (anonymous):

so can you show me?

OpenStudy (anonymous):

\[g(g^{-1}(x)) =x \]\[\therefore\ g(g^{-1}(x)) = 3(g^{-1}(x)) - 5\]\[\implies x = 3(g^{-1}(x)) - 5\] Now solve for \(g^{-1}(x)\)

OpenStudy (anonymous):

so what is g

OpenStudy (anonymous):

is it going to equal x + 5 /3

OpenStudy (anonymous):

or one of these 3x + 5 5 - 3x -3x + 5

OpenStudy (anonymous):

Did you solve for it?

OpenStudy (anonymous):

Add 5 to both sides,the divide both sides by 3.

OpenStudy (anonymous):

so its going to be x+5/3